Isolation of eugenol lab by extraction and simple distillation

20ml of dichloromethane (1g clove oil / 10ml dichloromethane) solution

I'm supposed to calculate the percent recovery of Eugenol in this mixture . However imnnot sure how I'm supposed to go on about it.
Initial weight of flask(empty) was 39.8g
Final weight of flask (with eugenol) 40.4g

How am I supposed to calculate the percent of eugenol obtained if I don't know how much clove oil I obtained as well.
Am I supposed to ovtain a percent recovery for eugenol too? If so how do I do it.

Thankyouuu
Answer #1

Eugenol is a vanilloid and one of the primary components of clove oil. It can be extracted from ground-up cloves using steam distillation. It has been reported that clove oil is approximately 17 % clove by weight of cloves.

Here, we have a dichloromethane solution where 10 mL of the solution has 1 g of clove oil.Thus, 20 mL of this solution will have 2 g of clove oil.

Therefore, the initial amount of cloves used is

0.17*w_{cloves}=2\Rightarrow w_{cloves}=11.76g (1)

where, wcloves is the initial weight og ground cloves used in the experiment.

We can calculate the weight of eugenol as

w_{eugenol}=w_{fl+eugenol}-w_{fl}=40.4-39.8=0.6g   (2)

where, wfl is the weight of empty flask, wfl+eugenol is the weight of the flask with eugenol and weugenol is the weight of eugenol obtained from cloves.

Thus, percentage recovery of Eugenol will be

\%recovery=\frac{w_{eugenol}}{w_{cloves}}*100\%=\frac{0.6g}{11.76g}*100\%=5.1\% (3)

and the percentage yield from the clove oil mixture would be

\%yield=\frac{w_{eugenol}}{w_{clove oil}}*100\%=\frac{0.6g}{2g}*100\%=30\% (4)

where, wcloveoil is the weight of clove oil in 20 mL of dichloromethane solution.